subject: operations research

When customers arrive at Cool's Ice Cream Shop, they take a number and wait to be called to purchase ice cream from one of the counter servers. From experience in past summers, the store's staff knows that customers arrive at a rate of 150 per hour on summer days between 3:00 p.m. and 10:00 p.m., and a server can serve 1 customer in 1 minute on average. Cool's wants to make sure that customers wait no longer than 10 minutes for service. Cool's is contemplating keeping three servers behind the ice cream counter during the peak summer hours.

(i)        Will this number be adequate to meet the waiting time policy?

(ii)       What will be the probability that 3 to 4 customers in Shop?

(iii)      In winter season, arrival rate of customer is reduced to half from 3:00 p.m. and 10:00 p.m. What decision should be taken by the owner according to cost cutting point of view?

Question 4b:

Analysis of arrivals at a PSO gas station with a single pump (filler) has shown the time between arrivals with a mean of 10 minutes. Service times were observed with a mean time of 6 minutes.

(i)        What is the probability that a car will have to wait?

(ii)       What is the mean number of customers at the station?

(iii)      What is the mean number of customers waiting to be served?

(iv)      PSO is willing to install a second pump when convinced that an arrival would expect to wait at least twelve minutes for the gas. By how much the flow of arrivals is increased in order to justify a second booth?                                      

please chow the calculations and formula.
Answer #1

As per Chegg guidelines, asan expert i have to answer only one question otherwise Chegg management will take action on my account. But i explained clearly two questions given by you. Please understand and Please up vote this answer. Tank you so much.

Answer 4a :

solution :-) rate of 150 per howy () Customer worrive at on Summer day between 3:00 pm and 10500PM. ora 2=150 => Server can s.: Se ven : pere howus.is 3x60 = 180 oo Rate of service U = 180 per hows: => Cools want to make sure that customers wait no lhave any -> hope you understand. Dt you doubt then comment it. Thank you

Answer 4b :

Average arrival rate, λ = 1 per 10 minutes = 6 per 60 minutes = 6 per hr.
Average service rate, μ = 1 in 6 minutes = 10 per 60 minutes = 10 per hr.

(i)

The probability of an arriving car waits = 1 - λ/μ = 1 - 6/10 = 0.4 or 40%

(ii)

The average number of customers in the station, Ls = λ / (μ - λ) = 6 / (10 - 6) = 1.50

(iii)

The average number of customers waiting, Lq = λ2 / {μ.(μ - λ)} = 6^2/(10*(10-6)) = 0.9 customrs

(iv)

The required average waiting time (Wq) is 12 minutes or 0.20 hours.

So,

Wq = λ / {μ.(μ - λ)} >= 0.2
or, λ / {10*(10 - λ)} >= 0.2
or, λ = 20 - 2λ
or, λ = 6.667

So, the arrival rate has to be 6.667 per hour (which is 1 customer in 60/6.667 i.e. 9 minutes) or more to justify adding another server.

__________________________________________________________________________________________________________

Please Please up vote this answer. Your up vote is so valuable to me. Thank you soo much.